# Zero and Nonzero Energy Majorana Modes in an Extended Kitaev Chain

### v1

@@ V Conclusion
In conclusion, we have investigated an extended Kitaev chain that contains three sites per unit cell, resulting from the hybridization of a trimer SSH model [50] with the standard Kitaev chain [1], which leads to a total of six Majorana sites per unit cell. This hexamer lattice provides a richer edge state structure, which manifests itself compared to existing extended Kitaev models [52, 53, 54, 55, 56]. The system is found to preserve chiral, time-reversal, and particle-hole symmetries. To analyse the role of these symmetries in protecting the edge states, a set of generic perturbations has been applied, some of which preserve all symmetries, whereas the others break some of the symmetries. The edge modes’ resilience has been confirmed by this investigation. Furthermore, we considered the effect of spatial disorder on the system. In particular, we apply the disorder individually to each of the system parameters, including the previously introduced perturbations. As a result, we find that the system demonstrates the expected topological protection across all cases, maintaining robust edge modes.
Even though the Kitaev chain has been thoroughly studied and successfully implemented on a variety of physical platforms [38, 39, 40, 41, 42, 43, 44, 45], it remains a fundamental concept that inspires a considerable range of ongoing research. We expect that this extended version of the Kitaev chain will garner substantial attention in both theoretical and experimental contexts. From the experimental perspective, the inclusion of both robust MZMs and Majorana nonzero energy edge modes in our model provides richer experimental signatures. This coexistence of different edge modes can improve detection techniques by offering more diverse experimental signals, thereby contributing to progress in the experimental identification of Majorana modes. On the theoretical side, the model presented in this paper is amenable to further modification for exploration under the lens of non-Hermitian systems [58, 59, 60, 61, 62, 63], Floquet systems [64, 65, 51, 66, 67, 68, 69, 70, 70, 71, 72], or interacting systems [73, 74]. Another promising future aspect is to consider replacing the trimerized SSH model [50] with one that includes more sites per unit cell and similarly combining it with the Kitaev chain [1], which we expect to further enrich the edge modes’ profiles and potentially lead to a different topological structure. Finally, the technique employed for constructing our model is likely to pique the researchers’ curiosity in exploring hybridizations beyond the SSH and Kitaev paradigms, which would give rise to other families of superconductors featuring enlarged unit cells and/or different topological characteristics.

### v2

@@ V Conclusion
In conclusion, we have investigated an extended Kitaev chain that contains three sites per unit cell, resulting from the hybridization of a trimer SSH model [30] with the standard Kitaev chain [36], which leads to a total of six Majorana sites per unit cell. This hexamer lattice provides a richer edge state structure, which manifests itself compared to existing extended Kitaev models [74, 2, 62, 52, 18, 76, 78]. The system is found to preserve chiral, time-reversal, and particle-hole symmetries. To analyse the role of these symmetries in protecting the edge states, a set of generic perturbations has been applied, some of which preserve all symmetries, whereas the others break some of the symmetries. The edge modes’ resilience has been confirmed by this investigation. Furthermore, we considered the effect of spatial disorder on the system. In particular, we apply the disorder individually to each of the system parameters, including the previously introduced perturbations. As a result, we find that the system demonstrates the expected topological protection across all cases, maintaining robust edge modes.
Even though the Kitaev chain has been thoroughly studied and successfully implemented on a variety of physical platforms [72, 13, 71, 24, 32, 3, 67, 14], it remains a fundamental concept that inspires a considerable range of ongoing research. We expect that this extended version of the Kitaev chain will garner substantial attention in both theoretical and experimental contexts. From the experimental perspective, the inclusion of both robust MZMs and Majorana nonzero energy edge modes in our model provides richer experimental signatures. This coexistence of different edge modes can improve detection techniques by offering more diverse experimental signals, thereby contributing to progress in the experimental identification of Majorana modes.
On the theoretical side, the model presented in this paper is amenable to further modification for further exploration of exotic topological phases. For instance, a simplified model that supports more analytically tractable non-zero energy edge modes is worth considering, as it allows the application of the transfer matrix method of Refs. [20, 41] to fully characterize their topology. Moreover, the addition of non-Hermiticity [80, 63, 26, 28, 68, 48, 70], periodic driving [82, 12, 31, 15, 10, 81, 8, 7, 7, 9, 37, 34], and/or interaction [61, 85], is expected to further enrich the system.
Another promising future aspect is to consider replacing the trimerized SSH model [30] with one that includes more sites per unit cell and similarly combining it with the Kitaev chain [36], which we expect to further enrich the edge modes’ profiles and potentially lead to a different topological structure. Finally, the technique employed for constructing our model is likely to pique the researchers’ curiosity in exploring hybridizations beyond the SSH and Kitaev paradigms, which would give rise to other families of superconductors featuring enlarged unit cells and/or different topological characteristics.
@@ Acknowledgements.
Appendix A Appendix section
In Figs. 10 and 11 we show the wavefunction profiles of the edge modes in our system at $N=30$ unit cells and $N=300$ unit cells, respectively. In both figures, the number of peaks for each edge mode remains unchanged despite using a different number of lattice sites. In particular, there are two dominant peaks at each of the finite-energy edge modes, and this number of peaks is invariant and consistent with the profiles shown in the main text for $N=100$ unit cells (Fig. 5), which confirms their system-size-independent nature.
Figure 10: Wave function profiles associated with the system’s edge states for $N=30$ unit cells at $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, and $\mu=3.5$. Note that each subplot displays one left-localized edge state and one right-localized edge state.
Figure 11: Wave function profiles associated with the system’s edge states for $N=300$ unit cells at $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, and $\mu=3.5$. Note that each subplot displays one left-localized edge state and one right-localized edge state. Each subplot contains inset zooms highlighting the peaks at both edges.